# PhysicsCorrect: A Training-Free Approach for Stable Neural PDE Simulations

### v1

@@ B.5 Sensitivity Analysis: Error Structure and Magnitude Limits
The diagonal dashed line serves as a reference indicating no correction. Across all systems, our method significantly reduces the prediction error compared to the uncorrected baseline, particularly when the corruption is structured (Gaussian random field). The analysis reveals a fundamental distinction between linear and nonlinear systems. For the wave equation, being a linear PDE with exact Jacobian computation, the corrector maintains effectiveness across the entire tested error range, indicating that the PhysicsCorrect module has likely reached the discretization error floor. The linear nature of the governing equation ensures that our correction mechanism remains theoretically sound regardless of error magnitude because the PDE residual is linear to the prediction, where Equation 5 does not rely on the assumption of small relative error on the prediction for Taylor expansion.

### v2

@@ B.6 Sensitivity Analysis: Error Structure and Magnitude Limits
The diagonal dashed line serves as a reference indicating no correction. Across all systems, our method significantly reduces the prediction error compared to the uncorrected baseline, particularly when the corruption is structured (Gaussian random field). The analysis reveals a fundamental distinction between linear and nonlinear systems. For the wave equation, being a linear PDE with exact Jacobian computation, the corrector maintains effectiveness across the entire tested error range, indicating that the PhysicsCorrect module has likely reached the discretization error floor. The linear nature of the governing equation ensures that our correction mechanism remains theoretically sound regardless of error magnitude because the PDE residual is linear to the prediction, where Equation 3.1 does not rely on the assumption of small relative error on the prediction for Taylor expansion.
@@
B.7 Sensitivity Analysis: Effect of Spatial Discretization and Resolution
Beyond the sensitivity analysis of prediction accuracy to correction performance, as discussed in the main text, the accuracy of PDE residual evaluation also plays a key role in determining correction quality. Inaccurate residuals can degrade the overall performance of the prediction–correction pipeline. Both the spatial discretization scheme and the spatial resolution of the physical field strongly influence the accuracy of the residual calculation. In this subsection, we further examine how discretization and resolution in the prediction step affect the performance of the prediction–correction framework.
Experimental Protocol.
[…]
Effect of Discretization.
As shown in Figure 15 (left), higher‑accuracy discretization schemes consistently yield better prediction–correction performance, as each correction step produces a more accurate corrected state and keeps the predicted trajectory closer to the reference solution. While advanced discretization schemes appear similar at a glance, we observe small but non‑negligible differences, with average discrepancies below $3.5\times 10^{-5}$. These results indicate that improving the spatial discretization enhances one‑step correction accuracy, which in turn benefits long‑term stability.
Effect of Resolution.
Similarly, Figure 15 (right) shows that reducing the spatial resolution of both the prediction and the reference field makes PDE residual evaluation more challenging, leading to larger errors in subsequent correction steps. Coarser grids exacerbate error accumulation, ultimately degrading prediction–correction performance. The coarser the grid, the worse the accuracy. We note that this limitation could be mitigated by integrating spatio-temporal decomposition training frameworks (Huang et al. 2023), preserving the large spatial- and temporal-interval advantages of purely neural PDE solvers while enhancing stability through physics-based correction.
Practical Implications.
The results collectively demonstrate that both spatial discretization quality and spatial resolution are critical for achieving stable and accurate prediction–correction performance. High‑accuracy discretization schemes improve the fidelity of PDE residual estimation, leading to more effective corrections and slower long‑term error growth. Likewise, higher spatial resolution mitigates residual evaluation errors and reduces the risk of error amplification during rollout. In practice, these two factors are intertwined: a coarse resolution can undermine the benefits of an advanced discretization scheme, whereas a high‑accuracy discretization can partly compensate for moderate resolution. Therefore, optimal performance is achieved when both discretization and resolution are chosen to balance computational cost with the accuracy requirements of the task.
[…]
B.8 Potential scalability improvements
In order to improve the scalability of the proposed approach, we conducted preliminary blockwise Jacobian–vector product experiments, which show a 34 % reduction in memory and runtime with comparable accuracy at 1282 resolution. The bottleneck stems from pseudoinverses rather than the correction principle. In addition, iterative CG solvers or structure-aware Jacobians would enable 3-D extension of our approach.
Appendix C Broader Impacts
Time-dependent PDEs are fundamental to scientific and engineering applications including fluid dynamics, material modeling, seismic analysis, and climate forecasting. While neural PDE solvers offer computational efficiency, their long-term stability remains challenging for reliable predictions. Our physics-informed correction framework addresses this limitation without requiring model retraining, enabling seamless integration with existing neural operators across diverse problems. This training-free approach particularly benefits high-stakes domains like climate modeling and autonomous systems, where error accumulation can compromise prediction reliability.
However, surrogate modeling carries inherent responsibilities. Our method’s effectiveness depends critically on accurate PDE residual formulation and evaluation. We emphasize that corrected predictions require careful verification, especially in applications where simulation outputs inform critical decisions. Users should validate results against known benchmarks and consider the method’s limitations, particularly regarding computational scaling and discretization errors, when deploying in production environments.